# Mergers of compact objects with cores of massive stars: evolution, r-process and multi-messenger signatures

### v1

@@ 5.1.5 Long gamma ray bursts
The mergers of BHs with cores of giant stars have been previously proposed as possible progenitors of LGRBs (e.g., Fryer & Woosley 1998; Zhang & Fryer 2001). The rapid accretion of matter from the rotating core through an accretion disk formed around the BH leads to launching of relativistic jets that could energize the LGRB explosions. The gamma-rays are produced by synchrotron radiation that is emitted as the jets interact with their surroundings, as in the case of collapsars that come from single stars (e.g., MacFadyen & Woosley 1999). We note, however, that when the merger occurs inside the opaque envelope of the giant the jets are likely to get chocked by the envelope gas (e.g., Papish et al. 2015; Grichener & Soker 2021) and trap the gamma-rays. Therefore, it is more likely that such LGRBs would be produced in delayed mergers between the naked core and the compact object, due to interaction of the binary with a circumbinary disk similarly to the scenario discussed for LFBOTs in section 5.1.4, or as a result of induced gravitational collapse of the NS to a BH due to accretion of the ejecta mass after the core explodes in a stripped-envelope SN event (e.g., Rueda & Ruffini 2012; Fryer et al. 2014). Since the envelope is ejected prior to the explosion(s), such LGRBs will not show hydrogen features in their lightcurves, and their exclusive association with SN Ic-BL (e.g., Drout et al. 2011; Modjaz et al. 2016; Japelj et al. 2018) might point towards mergers that occur past helium depletion in the core (e.g., Dessart et al. 2017).
@@
5.2 High energy neutrinos messengers
@@
The propagation of the collimation and internal shocks in the jet can accelerate protons to very high energies through the diffusive shock acceleration mechanism, also known as first-order Fermi acceleration (e.g., Bell 1978; Blandford & Ostriker 1978). The protons repeatedly cross the front of the shock due to scattering caused by magnetic fields irregularities. Since the typical timescale for this scattering is much shorter than the timescales for particle collisions in the shock, i.e, the shocks can be regarded as collisionless, the protons can gain energy at the expense of the difference in the kinetic energy between the two sides of the shock. Moreover, Grichener & Soker (2021) showed that for the typical luminosities of jets launched inside the CE, the upstream shock is optically thin to electron scattering and to pair production, preventing declaration of the accelerated protons by interactions with lower energy electrons that can be produced through these processes 666Guarini et al. (2023) study particle acceleration in magnetically driven jets and concludes that if these jets are halted in an extended envelope no efficient particle acceleration can occur at the collimation shocks due to high optical depths, but they do not rule out particle acceleration in internal shocks formed in an hydrodynamical jets chocked in an extended envelope. . Jets produced during the merger itself, however, are much more luminous, and therefore their shocks would be mediated by radiation, preventing efficient acceleration of protons and emission of very energetic neutrinos.
@@
While the energy flux of high energy neutrinos and gamma-rays should be comparable, the isotropic diffused gamma-rays emission measured by Fermi gamma-ray telescope (Ackermann et al., 2015) is much lower than the isotropic neutrino flux inferred by IceCube (e.g., Aartsen et al. 2014), pointing towards an optically thick high energy neutrinos source. Grichener & Soker (2021) showed that the jets launched by a BH while spiraling inside the envelope of a giant are chocked inside the extended envelope, which is optically thick to gamma-rays in the sub-TeV range, in which the gamma-rays that accompany high energy neutrino emission are expected to be detected due to cooling by interaction with matter (e.g., Celli et al. 2017; Capanema et al. 2021).
@@
The production of high energy neutrinos during mass accretion by a stellar mass BH that resides in a binary system has been studied in the literature in other scenarios as well. Sridhar et al. (2024), for instance, study the emission of high energy neutrinos from winds and jets launched by a compact object that accretes mass from an evolved post-MS companion at a super-Eddington rate following RLOF but shortly prior to CEE. Such accretion can power an hypernebula (Sridhar & Metzger, 2022), which is an energetic bubble () inflated into the circumbinary medium and is though to be correlated with recurring fast radio bursts (e.g., Sridhar et al. 2021). The neutrinos are mostly produced through photohadronic interactions of protons that are accelerated at the termination shocks with photons from the accretion disk. They find that hyperneubulae could explain up to of the high energy diffuse neutrino flux inferred by IceCube.
@@
Guarini et al. (2022) study high energy neutrinos emission during the merger of a BH with the stripped core of a giant after a stable RLOF, in a scenario that could result in LFBOTs (Metzger 2022; see section 5.1.4). The high mass accretion rates lead to the formation of an accretion disk that ejects outflows in the polar regions. These outflows propagate in the CSM generating shocks where protons can be accelerated. The protons then interact with steady target protons of the CSM, producing high energy neutrinos as they cool. The neutrino flux emitted in this scenario is expected to be below the sensitivity of IceCube. Future detection of neutrinos correlated with LFBOTs would be of crucial importance both for probing the origin of high energy neutrinos, and shedding light on progenitors of LFBOTs and perhaps on the physics of BH-core mergers.
5.3 Gravitational waves messengers
@@
Another multi-messenger signature that could be invaluable for associating NS/BH-core mergers with transients is GWs. Over the past years, several studies have been done to explore GWs emission from the in-spiral during the CEE phase and its detectabllity prospects with current and near future facilitates. Renzo et al. (2021) perform a model free study that focuses on GWs emission during the self-regulated phase of CEE and find it is potentially possible to expect at least a few GW detections from Galactic CEE within the LISA mission (e.g., Thorpe et al. 2019). Morán-Fraile et al. (2023) compute the GWs signals from the dynamical CEE phase of low mass stars using three dimensional magneto-hydrodynamics simulations and find that the final stages of CEE leading to the merger of the stellar cores are in the frequency band of LISA observations.
@@
Holgado et al. (2018) and Ginat et al. (2020) study GWs signatures from a CEE phase of an NS with a massive star. While Holgado et al. (2018) focus on the in-spiral and conclude that it might result in GW emission detectable by advanced LIGO as far away as the Magellanic Clouds, with stronger signals than low mass X-ray binaries, Ginat et al. (2020) explicitly study GWs emission during the merger of the NS with the core. They find that the in-spiral leading to such merger events produces a GW signature that differs significantly from the in-spiral leading to binary compact objects merger. This signature could be observable by space-based GW detectors, but due to the short duration of the event they expect it is likely to be detected with BBO (e.g., Cutler & Holz 2009) rather than LISA. We note, however, that since Ginat et al. 2020 do not simulate the feedback of the in-spiral on the surrounding gas, their results might differ from the expected observations.
@@
Another source of potentially detectable GWs produced during the merger of a compact object with the core of a giant star are Rossby instabilities that form in the accretion disk as a result of sharp density gradients. Non-axisymmetric modes of these instabilities could induce density waves, creating a quadruple moment that generates GWs. Gottlieb et al. (2024) explore the GWs emission from collpsars disks and find they may be detectable in LIGO-Virgo-Kagra (e.g., Abbott et al. 2018; Abbott et al. 2020) at a rate of about one event per year, implying that existing data might already contain such GW events. Turbulent motions in the jet-cocoon outflow could also result in stochastic GW bursts with a wide range of frequencies. Gottlieb et al. (2023) study the GW bursts powered by cocoons in collapsars and find they emit in the LIGO-Virgo-Kagra band and could be detected with third generation facilities like Cosmic Explorer (e.g., Evans et al. 2023) and the Einstein Telescope (e.g., Punturo et al. 2010).
@@
Understating the detectability prospects of these non-spiralling GW sources in the case of NS/BH-core mergers requires to combine an estimate for their event rates (e.g., Schrøder et al. 2020; Grichener 2023) with the expected frequencies and amplitudes of the GW signal. While NS/BH-core mergers are expected to occur more frequently than collpsars, it is challenging to predict the difference in their GW signature from analytical considerations. Computing the GWs patterns that originate in the accretion disk and cocoons formed around the jets during NS/BH-core merger events requires a treatment in the framework of full three dimensional hydrodynamical simulations of these systems due to the asymmetric nature of jets launching. Identifying GWs signals that are compatible with predictions of these simulations could be of crucial importance in shedding light on the physics of NS/BH-core mergers and on the jet launching mechanism, and advance our understanding of these transients.
@@ 6 How do we move forward?
Conducting three dimensional hydrodynamical simulations of NS/BH-core mergers is essential for accurately predicting their expected observational signatures and confidently linking them to detected transients. The subject of mass accretion by a compact object as it spirals trough envelope during the CEE phase has been broadly studied using hydrodynamical simulations (see Röpke & De Marco 2023 for a recent review). Though several groups studied the evolutionary stages prior to the merger (e.g., Schrøder et al. 2020) or up the formation of an accretion disk (e.g., Everson et al. 2024), to date, there are no three dimensional hydrodynamical simulations that capture jets launching and evolution during NS/BH-core mergers. Performing hydrodynamical simulations of such systems can be very challenging computationally. One of the most problematic numerical difficulties arises from the fact that it contains two processes that occur on very different scales. The Keplerian orbital motion of the compact object is extremely slow compared to the velocity of the sub-relativistic or relativistic jets that it launches. This creates a problem in defining a typical timestep suitable for running the simulation, since setting small enough timesteps as required to resolve the launching and interactions of mildly relativistic jets prevents us from simulating even a considerable fraction of one orbit of the compact object. López-Cámara et al. (2020b), for instance, inject the jets that an NS launches in the envelope of a giant star with a velocity of and are able to simulate only a small fraction of the NS’s orbit. Even without accounting for the orbital motion, resolving the launching of jets by a compact object can be computationally challenging due to the small scales involved (see section 3 in Soker 2024 for a discussion). One way to address this challenge is by using ’subgrid physics’ to mimic the effect of jets launching in the simulations, i.e., by approximating the collision of the jets’ material with its immediate surroundings (e.g., Schreier et al. 2021; Hillel et al. 2022a; Schreier et al. 2023; Hillel et al. 2023). Other complications arise from the variety of physical ingredients that are important for simulating such mergers, such as the self-gravity of the giant star, the magnetic fields prior to the jets launching and in the jets, relativistic effects related to the jets propagation in the envelope or core media, neutrino transport and the effects of radiation pressure and neutrino cooling on the accretion rate by the compact object (see section 3 for more details). Advances in computational power may enable the execution of such simulations in the future, but even partial simulations that omit some physical processes can still offer invaluable insights for their detection.
@@
Performing three dimensional hydrodynamical simulations of accreting compact objects inside the envelope and core of a giant star will advance research in this field and open new research directions that can lead to a better understanding of these systems and their respective outcomes. Post-processing of these simulations is the next crucial step to explore the possible processes that can occur inside the accretion disk and the jets it launches. To find the contribution of NS-core mergers to r-process abundances in the Galaxy and Universe (section 4) we can use the properties of the accretion disk and jets material obtained in these simulations, with emphasis on their neutron richness, and insert them to a nuclear reaction network like skynet (Lippuner & Roberts, 2017) to compute the r-process nucleosynthesis yields in such events; The velocity profiles and nucleosynthesis yields from the merger could be given as an input to a radiation hydrodynamic code to calculate the expected lightcurves (section 5.1); Computing the estimated flux of high energy neutrinos (section 5.2) emitted prior to a BH-core merger event and giving detectability prospects with IceCube (e.g., Aartsen et al. 2013) or IceCube-GEN2 (e.g., Clark et al. 2021) could be done by obtaining the energies of the protons accelerated across the shocks interface and taking into account all their possible cooling processes as well as the detailed micro-physics of pion production and decay; The GWs signatures and the potential to identify them in the O4 data or with future runs/next generation ground or space-based detectors could be obtained from using retarded Green’s functions to post-process the simulations output (section 5.3; see the Appendix in Gottlieb et al. 2023 for an implementation). A simultaneous detection of multiple messengers from the same transient, aligning with predictions from hydrodynamic simulations, could serve as a ’smoking gun’ to confirm its origin as an NS/BH-core merger.
@@ 7 Summary
Evolution. Massive binaries can evolve into NS/BH-core mergers if the SN explosion in which the primary star produces a compact object does not disrupt the binary system. In such cases the secondary star may engulf the compact object, initiating a CEE phase. Depending on the properties of the compact object-giant binary system at the onset of CEE, it could lead to either the ejection of the shared envelope and formation of a binary compact objects system, or the merger of the compact object with the core of the giant. The merger could occur either trough tidal disruption of the core’s matter by the compact object before it manages to reach the core’s surface, or through disruption of the core by the vast accretion that occurs when the compact object spirals inside it. In both cases the accretion disk that is formed around the compact object could launch jets that power a luminous transient.
R-process. If the rate at which an NS accretes mass as it merges with the core of a giant star is high enough, the hot and dense matter of the neutrino-cooled accretion disk that is formed around the NS can self neutronize and produce lanthanides through captures of neutrons on isotopes that originate in cooler regions of the disk. This r-process nucleosynthesis occurs on timescales that are dictated by the evolution of the secondary star ( tens Myrs), potentially allowing NS-core mergers to account for heavy elements formation in the early Universe and UFD galaxies, where BNS mergers face challenges. Three dimensional hydro-dynamical simulations are required to properly simulate r-process nucleosynthesis in NS-core mergers.
Multi-messenger signatures. The disk winds and jets launched by the accretion disk formed around the compact object during the merger play a key role in shaping the lightcurve of the resultant transient. They lead to the formation of a highly asymmetric CSM, contribute to the variability of the emission and produce dust with strong IR radiation. The mergers of NSs/BHs with cores of giant stars have been proposed as progenitors of transients that at least partially exhibit these behaviours, such as type IIn SNe, iPTF14hls-like transients, LFBOTs and low-luminosity LGRBs. However, since the gas of the giant’s envelope is very opaque, not allowing for direct optical observations of the underlying physical processes, detecting GWs and high energy neutrinos signatures correlated with these mergers would be of crucial importance for associating a transient with a merger origin.

### v2

@@ 5.1.5 Long gamma ray bursts
‘ The mergers of BHs with cores of giant stars have been previously proposed as possible progenitors of LGRBs (e.g., Fryer & Woosley 1998; Zhang & Fryer 2001). The rapid accretion of matter from the rotating core through an accretion disk formed around the BH leads to launching of relativistic jets that could energize the LGRB explosions. The gamma-rays are produced by synchrotron radiation that is emitted as the jets interact with their surroundings, as in the case of collapsars that come from single stars (e.g., MacFadyen & Woosley 1999). We note, however, that when the merger occurs inside the opaque envelope of the giant the jets are likely to get choked by the envelope gas (e.g., Papish et al. 2015; Grichener & Soker 2021) and trap the gamma-rays. Therefore, it is more likely that such LGRBs would be produced in delayed mergers between the naked core and the compact object, due to interaction of the binary with a circumbinary disk similarly to the scenario discussed for LFBOTs in section 5.1.4, or as a result of induced gravitational collapse of the NS to a BH due to accretion of the ejecta mass after the core explodes in a stripped-envelope SN event (e.g., Rueda & Ruffini 2012; Fryer et al. 2014). Since the envelope is ejected prior to the explosion(s), such LGRBs will not show hydrogen features in their lightcurves, and their exclusive association with SN Ic-BL (e.g., Drout et al. 2011; Modjaz et al. 2016; Japelj et al. 2018; Finneran et al. 2024) might point towards mergers that occur past helium depletion in the core (e.g., Dessart et al. 2017).
@@
5.2 High energy neutrino messengers
@@
The propagation of the collimation and internal shocks in the jet can accelerate protons to very high energies through the diffusive shock acceleration mechanism, also known as first-order Fermi acceleration (e.g., Bell 1978; Blandford & Ostriker 1978). The protons repeatedly cross the front of the shock due to scattering caused by magnetic field irregularities. Since the typical timescale for this scattering is much shorter than the timescales for particle collisions in the shock, i.e, the shocks can be regarded as collisionless, the protons can gain energy at the expense of the difference in the kinetic energy between the two sides of the shock. Moreover, Grichener & Soker (2021) showed that for the typical luminosities of jets launched inside the CE, the unshocked medium ahead of the shock, known as the upstream of the shock, is optically thin to electron scattering and to pair production, preventing deceleration of the accelerated protons by interactions with lower energy electrons that can be produced through these processes 666Guarini et al. (2023) study particle acceleration in magnetically driven jets and concludes that if these jets are halted in an extended envelope no efficient particle acceleration can occur at the collimation shocks due to high optical depths, but they do not rule out particle acceleration in internal shocks formed in an hydrodynamical jets choked in an extended envelope. . Jets produced during the merger itself, however, are much more luminous, and therefore their shocks would be mediated by radiation, preventing efficient acceleration of protons and emission of very energetic neutrinos.
@@
While the energy flux of high energy neutrinos and gamma-rays should be comparable, the isotropic diffused gamma-rays emission measured by Fermi gamma-ray telescope (Ackermann et al., 2015) is much lower than the isotropic neutrino flux inferred by IceCube (e.g., Aartsen et al. 2014), pointing towards an optically thick high energy neutrino source. Grichener & Soker (2021) showed that the jets launched by a BH while spiraling inside the envelope of a giant are choked inside the extended envelope, which is optically thick to gamma-rays in the sub-TeV range, in which the gamma-rays that accompany high energy neutrino emission are expected to be detected due to cooling by interaction with matter (e.g., Celli et al. 2017; Capanema et al. 2021).
@@
The production of high energy neutrinos during mass accretion by a stellar mass BH that resides in a binary system has been studied in the literature in other scenarios as well. Sridhar et al. (2024), for instance, study the emission of high energy neutrinos from winds and jets launched by a compact object that accretes mass from an evolved post-MS companion at a super-Eddington rate following RLOF but shortly prior to CEE. Such accretion can power an hypernebula (Sridhar & Metzger, 2022), which is an energetic bubble () inflated into the circumbinary medium and is thought to be correlated with recurring fast radio bursts (e.g., Sridhar et al. 2021). The neutrinos are mostly produced through photohadronic interactions of protons that are accelerated at the termination shocks with photons from the accretion disk. They find that hypernebulae could explain up to of the high energy diffuse neutrino flux inferred by IceCube.
@@
Guarini et al. (2022) study high energy neutrino emission during the merger of a BH with the stripped core of a giant after a stable RLOF, in a scenario that could result in LFBOTs (Metzger 2022; see section 5.1.4). The high mass accretion rates lead to the formation of an accretion disk that ejects outflows in the polar regions. These outflows propagate in the CSM generating shocks where protons can be accelerated. The protons then interact with steady target protons of the CSM, producing high energy neutrinos as they cool. The neutrino flux emitted in this scenario is expected to be below the sensitivity of IceCube. Future detection of neutrinos correlated with LFBOTs would be of crucial importance both for probing the origin of high energy neutrinos, and shedding light on progenitors of LFBOTs and perhaps on the physics of BH-core mergers.
5.3 Gravitational wave messengers
@@
Another multi-messenger signature that could be invaluable for associating NS/BH-core mergers with transients is GWs. Over the past years, several studies have been done to explore GW emission from the in-spiral during the CEE phase and its detectabllity prospects with current and near future facilitates. Renzo et al. (2021) perform a model free study that focuses on GW emission during the self-regulated phase of CEE and find it is potentially possible to expect at least a few GW detections from Galactic CEE within the LISA mission (e.g., Thorpe et al. 2019). Morán-Fraile et al. (2023) compute the GW signals from the dynamical CEE phase of low mass stars using three dimensional magneto-hydrodynamics simulations and find that the final stages of CEE leading to the merger of the stellar cores are in the frequency band of LISA observations.
@@
Holgado et al. (2018) and Ginat et al. (2020) study GW signatures from a CEE phase of an NS with a massive star. While Holgado et al. (2018) focus on the in-spiral and conclude that it might result in GW emission detectable by advanced LIGO as far away as the Magellanic Clouds, with stronger signals than low mass X-ray binaries, Ginat et al. (2020) explicitly study GW emission during the merger of the NS with the core. They find that the in-spiral leading to such merger events produces a GW signature that differs significantly from the in-spiral leading to binary compact objects merger. This signature could be observable by space-based GW detectors, but due to the short duration of the event they expect it is likely to be detected with BBO (e.g., Cutler & Holz 2009) rather than LISA. We note, however, that since Ginat et al. 2020 do not simulate the feedback of the in-spiral on the surrounding gas, their results might differ from the expected observations.
@@
Another source of potentially detectable GWs produced during the merger of a compact object with the core of a giant star are Rossby instabilities that form in the accretion disk as a result of sharp density gradients. Non-axisymmetric modes of these instabilities could induce density waves, creating a quadruple moment that generates GWs. Gottlieb et al. (2024) explore the GW emission from collpsars disks and find they may be detectable in LIGO-Virgo-Kagra (e.g., Abbott et al. 2018; Abbott et al. 2020) at a rate of about one event per year, implying that existing data might already contain such GW events. Turbulent motions in the jet-cocoon outflow could also result in stochastic GW bursts with a wide range of frequencies. Gottlieb et al. (2023) study the GW bursts powered by cocoons in collapsars and find they emit in the LIGO-Virgo-Kagra band and could be detected with third generation facilities like Cosmic Explorer (e.g., Evans et al. 2023) and the Einstein Telescope (e.g., Punturo et al. 2010).
@@
Understanding the detectability prospects of these non-spiralling GW sources in the case of NS/BH-core mergers requires to combine an estimate for their event rates (e.g., Schrøder et al. 2020; Grichener 2023) with the expected frequencies and amplitudes of the GW signal. While NS/BH-core mergers are expected to occur more frequently than collpsars, it is challenging to predict the difference in their GW signature from analytical considerations. Computing the GW patterns that originate in the accretion disk and cocoons formed around the jets during NS/BH-core merger events requires a treatment in the framework of full three dimensional hydrodynamical simulations of these systems due to the asymmetric nature of jet launching. Identifying GW signals that are compatible with predictions of these simulations could be of crucial importance in shedding light on the physics of NS/BH-core mergers and on the jet launching mechanism, and advance our understanding of these transients.
@@ 6 How do we move forward?
Modeling NS/BH-core mergers is crucial for advancing towards these associations. Over the years, many studies performed one dimensional and two dimensional simulation mimicking the CEE phase involving a compact object and the core of a giant.(e.g., Taam et al. 1978b; Delgado 1980; Bodenheimer & Taam 1984; Armitage & Livio 2000; Gilkis et al. 2019; Grichener et al. 2021), or the merger itself (e.g., Soker et al. 2019; Grichener & Soker 2019; Grichener & Soker 2023). Performing a wider parameter space exploration with one dimensional simulations could give insight into their typical timescales, nucleosynthesis and locations within their host galaxies. Two dimensional simulations could also be used to study different aspects of the accretion flow and disk formation, as done in previous works (e.g., Shima et al. 1985; Fryer et al. 1996; Blondin & Pope 2009; Blondin 2013; Tripathi et al. 2024). Due to the highly asymmetrical nature of jets physics, however, conducting three dimensional hydrodynamical simulations of NS/BH-core mergers is essential for accurately predicting their expected lightcurves and remnant morphologies, which are crucial for confidently linking them to detected transients.
@@
The subject of mass accretion by a compact object as it spirals trough envelope during the CEE phase has been broadly studied using three dimensional simulations (see Röpke & De Marco 2023 for a recent review). Though several groups studied the evolutionary stages prior to the merger (e.g., Schrøder et al. 2020) or up the formation of an accretion disk (e.g., Everson et al. 2024), to date, there are no three dimensional hydrodynamical simulations that capture jet launching and evolution during NS/BH-core mergers. Performing hydrodynamical simulations of such systems can be very challenging computationally. One of the most problematic numerical difficulties arises from the fact that it contains two processes that occur on very different scales. The Keplerian orbital motion of the compact object is extremely slow compared to the velocity of the sub-relativistic or relativistic jets that it launches. This creates a problem in defining a typical timestep suitable for running the simulation, since setting small enough timesteps as required to resolve the launching and interactions of mildly relativistic jets prevents us from simulating even a considerable fraction of one orbit of the compact object. López-Cámara et al. (2020b), for instance, inject the jets that an NS launches in the envelope of a giant star with a velocity of and are able to simulate only a small fraction of the NS’s orbit. Even without accounting for the orbital motion, resolving the launching of jets by a compact object can be computationally challenging due to the small scales involved (see section 3 in Soker 2024 for a discussion). One way to address this challenge is by using ’subgrid physics’ to mimic the effect of jet launching in the simulations, i.e., by approximating the collision of the jets’ material with its immediate surroundings (e.g., Schreier et al. 2021; Hillel et al. 2022a; Schreier et al. 2023; Hillel et al. 2023). Other complications arise from the variety of physical ingredients that are important for simulating such mergers, such as the self-gravity of the giant star, the magnetic fields prior to the jet launching and in the jets, relativistic effects related to the jet propagation in the envelope or core media, neutrino transport and the effects of radiation pressure and neutrino cooling on the accretion rate by the compact object (see section 3 for more details). Advances in computational power may enable the execution of such simulations in the future, but even partial simulations that omit some physical processes can still offer invaluable insights for their detection.
@@
Performing three dimensional hydrodynamical simulations of accreting compact objects inside the envelope and core of a giant star will advance research in this field and open new research directions that can lead to a better understanding of these systems and their respective outcomes. Post-processing of these simulations is the next crucial step to explore the possible processes that can occur inside the accretion disk and the jets it launches. To find the contribution of NS-core mergers to r-process abundances in the Galaxy and Universe (section 4) we can use the properties of the accretion disk and jet material obtained in these simulations, with emphasis on their neutron richness, and insert them to a nuclear reaction network like skynet (Lippuner & Roberts, 2017) to compute the r-process nucleosynthesis yields in such events; The velocity profiles and nucleosynthesis yields from the merger could be given as an input to a radiation hydrodynamic code to calculate the expected lightcurves (section 5.1); Computing the estimated flux of high energy neutrinos (section 5.2) emitted prior to a BH-core merger event and giving detectability prospects with IceCube (e.g., Aartsen et al. 2013) or IceCube-GEN2 (e.g., Clark et al. 2021) could be done by obtaining the energies of the protons accelerated across the shocks interface and taking into account all their possible cooling processes as well as the detailed micro-physics of pion production and decay; The GW signatures and the potential to identify them in the O4 data or with future runs/next generation ground or space-based detectors could be obtained from using retarded Green’s functions to post-process the simulations output (section 5.3; see the Appendix in Gottlieb et al. 2023 for an implementation). A simultaneous detection of multiple messengers from the same transient, aligning with predictions from hydrodynamic simulations, could serve as a ’smoking gun’ to confirm its origin as an NS/BH-core merger.
@@ 7 Summary
Evolutionary pathways. Massive binaries can evolve into NS/BH-core mergers if the SN explosion in which the primary star produces a compact object does not disrupt the binary system. In such cases the secondary star may engulf the compact object, initiating a CEE phase. Depending on the properties of the compact object-giant binary system at the onset of CEE, it could lead to either the ejection of the shared envelope and formation of a binary compact objects system, or the merger of the compact object with the core of the giant. The merger could occur either through tidal disruption of the core’s matter by the compact object before it manages to reach the core’s surface, or through disruption of the core by the vast accretion that occurs when the compact object spirals inside it. In both cases the accretion disk that is formed around the compact object could launch jets that power a luminous transient.
R-process nucleosynthesis. If the rate at which an NS accretes mass as it merges with the core of a giant star is high enough, the hot and dense matter of the neutrino-cooled accretion disk that is formed around the NS can self neutronize and produce lanthanides through captures of neutrons on isotopes that originate in cooler regions of the disk. This r-process nucleosynthesis occurs on timescales that are dictated by the evolution of the secondary star ( tens Myrs), potentially allowing NS-core mergers to account for heavy element formation in the early Universe and UFD galaxies, where BNS mergers face challenges. Three dimensional hydro-dynamical simulations are required to properly simulate r-process nucleosynthesis in NS-core mergers.
Multi-messenger signatures. The disk winds and jets launched by the accretion disk formed around the compact object during the merger play a key role in shaping the lightcurve of the resultant transient. They lead to the formation of a highly asymmetric CSM, contribute to the variability of the emission and produce dust with strong IR radiation. The mergers of NSs/BHs with cores of giant stars have been proposed as progenitors of transients that at least partially exhibit these behaviours, such as type IIn SNe, iPTF14hls-like transients, LFBOTs and low-luminosity LGRBs. However, since the gas of the giant’s envelope is very opaque, not allowing for direct optical observations of the underlying physical processes, detecting GW and high energy neutrino signatures correlated with these mergers would be of crucial importance for associating a transient with a merger origin.